Suppose S,T element of L(V) are such that ST = TS. Prove that null S is invariant under T.


1
Expert's answer
2021-06-17T19:10:13-0400

solution


for any "\upsilon" "\in" nul (S), S("\upsilon") = 0

Since ST=T(T("\upsilon"))=T(S("\upsilon")) =T(0),

Then T("\upsilon")"\in" nul (S) invariant under T

"\upsilon"